OpenStudy (anonymous):

what is the trig identity of cos^2x-1=

OpenStudy (unklerhaukus):

\[\sin ^2 x+\cos^2x=1\]\[\cos^2x-1=\]

OpenStudy (anonymous):

can u answer please i have a better question;